# Admissibility of Stein Shrinkage for Batch Normalization in the Presence of Adversarial Attacks

### v1

@@ 4.2 Application to Cityscapes Data
BN Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
@@
Our (8) 68.8 $\pm$ 0.1 39.6 $\pm$ 2.3 32.5 $\pm$ 2.6 29.9 $\pm$ 2.4 28.9 $\pm$ 2.3 27.2 $\pm$ 1.8
Our (16) 68.4 $\pm$ 0.4 36.3 $\pm$ 2.7 29.4 $\pm$ 1.5 27.1 $\pm$ 1.5 26.3 $\pm$ 1.7 25.1 $\pm$ 2.2
Khoshsiratkhoshsirat2023improvingnormalizationjamessteinestimator (8) 68.4 $\pm$ 0.3 36.4 $\pm$ 2.8 28.5 $\pm$ 3.5 25.9 $\pm$ 3.5 24.8 $\pm$ 3.2 23.3 $\pm$ 3
Khoshsiratkhoshsirat2023improvingnormalizationjamessteinestimator (16) 67.1 $\pm$ 0.3 36.2 $\pm$ 0.7 28.7 $\pm$ 1.3 26.3 $\pm$ 1.3 25.5 $\pm$ 1.2 24.4 $\pm$ 1.2
@@
Table 2: Performance of HRNetV2 with different BN methods under varying subgaussian noise and batch sizes respectively. Values depict mean of mIoU scores $\pm$ standard deviation.
To further evaluate the robustness of different BN variants, we tested them on the Cityscapes dataset—a more challenging task involving high-resolution urban scene segmentation with fine-grained object boundaries. The dataset contains 5,000 finely annotated images from 50 cities with 19 semantic classes used for evaluation. As in the previous experiment, we injected subgaussian noise into the feature maps during validation and testing to simulate corrupted statistics. This setup is particularly relevant for segmentation, where maintaining accurate local feature distributions is crucial for pixel-level predictions. As shown in Table 2, standard BN fails to retain meaningful performance even under mild noise (e.g., below $10$ mIoU score at $3\%$ corruption). In contrast, our proposed method and its mean corrected variant (see the appendix) significantly improve robustness, maintaining mIoU scores above $25\%$ even at high noise settings. Regularization-based approaches (Lasso and Ridge) show negligible improvement under noise and have a behavior similar to that of standard BN.
@@ 4.3 Application to PPMI Data
In this experiment, the goal is to demonstrate the performance of our BN method on 3D diffusion tensor magnetic resonance images (DT-MRI or in short DTI) acquired from Parkinson’s Disease (PD) affected brains and normal brains (Controls). DTI is a non-invasive MRI technique that allows one to infer neuronal connectivity patterns in the tissue being imaged. The PPMI data was subjected to a standard pre-processing pipeline (see Appendix B) using the FSL software FSLpack . DTIs are matrix-valued images where each voxel has a $(3,3)$ symmetric positive definite matrix that captures the covaraince of a local Gaussian fit. These DTIs are further processed to extract scalar-valued feature maps, called fractional anisotropy (FA) maps. The FA is a scalar function of the three eigen values of the diffusion tensor at each voxel in the DTI. Harmonization FORTIN2017149 is applied to these FA maps to remove site effects caused by distinct acquisition protocols and equipment used. The data set is highly imbalanced and contains FA maps from 360 Control and 130 PD brain scans. This imbalance is addressed using mixup-basedDBLP:journals/corr/abs-1710-09412 augmentation with sampling bias favoring the smaller sized class (PD). Using the same setup as in previous experiments and a 5-fold cross-validation to compare 3D CNN with different variants of 3D BN, we get results summarized in the table 3. These results further support our earlier observations, even in this challenging classification task.
BN Method 3% noise 10% noise 15% noise 20% noise 30% noise
Standard (10) 71.2 $\pm$ 1.6 64.7 $\pm$ 4.3 62.3 $\pm$ 5 61.3 $\pm$ 3.9 60.7 $\pm$ 8.2
Standard (16) 72.6 $\pm$ 1.5 63.7 $\pm$ 6.7 63.5 $\pm$ 4.6 62.3 $\pm$ 5.6 59.1 $\pm$ 5.4
Our (10) 71.8 $\pm$ 2.8 68 $\pm$ 3.6 68 $\pm$ 3.9 66.8 $\pm$ 5.2 66.5 $\pm$ 6.9
Our (16) 72.4 $\pm$ 2.8 65.7 $\pm$ 3.6 63.5 $\pm$ 6 61.5 $\pm$ 5.3 60.9 $\pm$ 7.2
Khoshsiratkhoshsirat2023improvingnormalizationjamessteinestimator (10) 71.4 $\pm$ 2.4 67.6 $\pm$ 4.5 66.4 $\pm$ 4.5 65.3 $\pm$ 7 64.9 $\pm$ 7.2
Khoshsiratkhoshsirat2023improvingnormalizationjamessteinestimator (16) 70.8 $\pm$ 1.9 67.2 $\pm$ 3.9 65.1 $\pm$ 6.1 64.3 $\pm$ 3.7 63.5 $\pm$ 4.6
Lasso (10) 64.9 $\pm$ 6.9 62.5 $\pm$ 9.4 62.3 $\pm$ 12.9 60.7 $\pm$ 9.3 57.5 $\pm$ 7.1
Lasso (16) 65.3 $\pm$ 3.9 64.5 $\pm$ 5.2 62.9 $\pm$ 4.9 59.1 $\pm$ 4.6 57.2 $\pm$ 4.4
Ridge (10) 68.8 $\pm$ 3.6 64.7 $\pm$ 4.1 64.5 $\pm$ 6.5 62.9 $\pm$ 4 61.7 $\pm$ 5.3
Ridge (16) 71 $\pm$ 1.8 64.3 $\pm$ 7.4 63.7 $\pm$ 3.3 62.1 $\pm$ 5.5 61.1 $\pm$ 2.6
@@
Table 3: 3D CNN performance with different BN methods under varying subgaussian noise and batch sizes respectively on PPMI data. Values depict mean accuracies $\pm$ standard deviation.
5 Conclusions
In this paper, we presented a novel theory involving the proof of dominance of the Stein’s shrinkage estimators of mean and variance of a Gaussian random variable in the presence of adversarial attacks. We applied Stein’s shirnkage to BN in a variety of deep networks and demonstrated superior performance, in comparison to other BN methods, on three benchmark data sets. One possible explanation for the improved performance is the connection between model performance under attack and the Lipschitz properties of the network DBLP:journals/corr/abs-2003-02460 . Since BN is known to affect the Lipschitz constant conf/nips/SanturkarTIM18 , it would be interesting to investigate the influence of our BN method on this aspect in future work.
6 Acknowledgments

### v2

@@ 4.2 Application to Cityscapes Data
Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
@@
Ours (8) 68.8 $\pm$ 0.1 39.6 $\pm$ 2.3 32.5 $\pm$ 2.6 29.9 $\pm$ 2.4 28.9 $\pm$ 2.3 27.2 $\pm$ 1.8
Ours (16) 68.4 $\pm$ 0.4 36.3 $\pm$ 2.7 29.4 $\pm$ 1.5 27.1 $\pm$ 1.5 26.3 $\pm$ 1.7 25.1 $\pm$ 2.2
Mean corrected (8) 67.80 $\pm$ 0.2 39.13 $\pm$ 2.6 31.8 $\pm$ 2.1 29.1 $\pm$ 2 28.1 $\pm$ 2 26.4 $\pm$ 2.2
Mean corrected (16) 68.1 $\pm$ 0.5 39.07 $\pm$ 1.7 32.8 $\pm$ 1 30.8 $\pm$ 1.2 30 $\pm$ 1.3 28.6 $\pm$ 1.2
Khoshsirat et al. (8) 68.4 $\pm$ 0.3 36.4 $\pm$ 2.8 28.5 $\pm$ 3.5 25.9 $\pm$ 3.5 24.8 $\pm$ 3.2 23.3 $\pm$ 3
Khoshsirat et al. (16) 67.1 $\pm$ 0.3 36.2 $\pm$ 0.7 28.7 $\pm$ 1.3 26.3 $\pm$ 1.3 25.5 $\pm$ 1.2 24.4 $\pm$ 1.2
@@
Table 2: Performance of HRNetV2 with different BN methods under varying subgaussian noise and batch sizes respectively. Values depict mean of mIoU scores $\pm$ standard deviation on part of validation Cityscapes data using as a test data. Numbers inside paranthesis in the far left column denote batch size.
To further evaluate the robustness of different BN variants, we tested them on the Cityscapes dataset – a more challenging task that involves high resolution urban scene segmentation with fine-grained object boundaries. The dataset contains 5,000 finely annotated images from 50 cities with 19 semantic classes used for evaluation. We use the mean intersection over union (mIoU), a standard metric used for semantic segmentation in the computer vision community. As in the previous experiment, we injected sub-gaussian noise into the feature maps during validation and testing to simulate corrupted statistics. This setup is particularly relevant for segmentation, where maintaining accurate local feature distributions is crucial for pixel-level predictions. As shown in Table 2, standard BN fails to retain meaningful performance even under mild noise (e.g., below $10$ mIoU score at $3\%$ corruption). In contrast, our proposed method and its mean corrected variant significantly improve robustness, maintaining mIoU scores above $25\%$ even at high noise settings. Regularization-based approaches (Lasso and Ridge) show negligible improvement under noise and have a behavior similar to that of standard BN.
@@ 4.3 Application to PPMI Data
In this experiment, the goal is to demonstrate the performance of our BN method on 3D diffusion tensor magnetic resonance images (DT-MRI or in short DTI) acquired from Parkinson’s Disease (PD) affected brains and normal brains (Controls). DTI is a noninvasive MRI technique that allows one to infer neuronal connectivity patterns in the tissue being imaged. The PPMI data were subjected to a standard preprocessing pipeline (see Appendix B). using the FSL software (Jenkinson et al., 2012). DTIs are matrix-valued images where each voxel has a $(3,3)$ symmetric positive definite matrix that captures the covaraince of a local (voxel-level) Gaussian fit. These DTIs are further processed to extract scalar-valued feature maps, called fractional anisotropy (FA) maps. The FA is a scalar function of the three eigen values of the diffusion tensor at each voxel in the DTI. Harmonization (Jean-Philippe et al., 2017) is applied to these FA maps to remove site effects caused by distinct acquisition protocols and equipment used. The data set is highly imbalanced and contains FA maps from 360 Control and 130 PD brain scans. This imbalance is addressed using mixup-based augmentation (Zhang et al., 2018) with sampling bias favoring the smaller sized class (PD). Using the same setup as in previous experiments and a 5-fold cross-validation to compare 3D CNN with different variants of 3D BN, we get results summarized in the Table 3. These results further support our earlier observations, even in this challenging classification task. Since the image data in this experiment are 3D volumetric images, the computational needs are much higher than when handling 2D image data sets, and we therefore used smaller batch sizes in the reported results. This helps reduce the computational footprint of the experiment.
Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
Standard (10) 94.9 $\pm$ 1.3 71.2 $\pm$ 1.6 64.7 $\pm$ 4.3 62.3 $\pm$ 5 61.3 $\pm$ 3.9 60.7 $\pm$ 8.2
Standard (16) 93.3 $\pm$ 1.5 72.6 $\pm$ 1.5 63.7 $\pm$ 6.7 63.5 $\pm$ 4.6 62.3 $\pm$ 5.6 59.1 $\pm$ 5.4
Ours (10) 95.1 $\pm$ 1.7 71.8 $\pm$ 2.8 68 $\pm$ 3.6 68 $\pm$ 3.9 66.8 $\pm$ 5.2 66.5 $\pm$ 6.9
Ours (16) 95.1 $\pm$ 1.8 72.4 $\pm$ 2.8 65.7 $\pm$ 3.6 63.5 $\pm$ 6 61.5 $\pm$ 5.3 60.9 $\pm$ 7.2
Mean corrected (10) 95.3 $\pm$ 2.6 68.6 $\pm$ 7.3 64.7 $\pm$ 6.1 64.1 $\pm$ 5.7 63.7 $\pm$ 6.1 62.3 $\pm$ 4.3
Mean corrected (16) 93.1 $\pm$ 2.4 69.2 $\pm$ 7.9 62.9 $\pm$ 12.2 61.1 $\pm$ 7.5 59.7 $\pm$ 10.9 58.7 $\pm$ 8.2
Khoshsirat et al. (10) 95.1 $\pm$ 2.4 71.4 $\pm$ 2.4 67.6 $\pm$ 4.5 66.4 $\pm$ 4.5 65.3 $\pm$ 7 64.9 $\pm$ 7.2
Khoshsirat et al. (16) 94.7 $\pm$ 2.5 70.8 $\pm$ 1.9 67.2 $\pm$ 3.9 65.1 $\pm$ 6.1 64.3 $\pm$ 3.7 63.5 $\pm$ 4.6
Lasso (10) 96.5 $\pm$ 3.0 64.9 $\pm$ 6.9 62.5 $\pm$ 9.4 62.3 $\pm$ 12.9 60.7 $\pm$ 9.3 57.5 $\pm$ 7.1
Lasso (16) 95.1 $\pm$ 1.9 65.3 $\pm$ 3.9 64.5 $\pm$ 5.2 62.9 $\pm$ 4.9 59.1 $\pm$ 4.6 57.2 $\pm$ 4.4
Ridge (10) 95.7 $\pm$ 1.7 68.8 $\pm$ 3.6 64.7 $\pm$ 4.1 64.5 $\pm$ 6.5 62.9 $\pm$ 4 61.7 $\pm$ 5.3
Ridge (16) 97.1 $\pm$ 1.3 71 $\pm$ 1.8 64.3 $\pm$ 7.4 63.7 $\pm$ 3.3 62.1 $\pm$ 5.5 61.1 $\pm$ 2.6
@@
Table 3: 3D CNN performance with different BN methods under varying subgaussian noise and batch sizes respectively on FA maps data. Values depict mean of accuracies $\pm$ standard deviation. Numbers inside paranthesis in the far left column denote batch size.
4.4 Experiments using Data Dependent Attacks
To validate our theoretical findings, we conducted a series of experiments using well-known adversarial attacks. While numerous attack methods have been proposed, our focus here is on confirming the theoretical insights rather than providing an extensive empirical comparison. Therefore, we restricted our study to two of the most widely used attacks, FGSM (Goodfellow et al., 2015) and it’s iterative version PGD. The results are presented below; for detailed experimental setups, we refer the reader to the appendix.
We first evaluate robustness against the FGSM attack. As before, we use a ResNet-9 model trained on the CIFAR-10 dataset and HRNetV2 model on the Cityscapes dataset, with the same hyperparameters as in the previous experiments. For consistency, we consider batch sizes matching those used in the earlier experiments and run each configuration five times. The results at different perturbation levels ($\epsilon=0.03,0.1,0.15,0.2,0.3$) are reported in Tables 4, 5. We omit regularized BN baselines on CIFAR-10, as they perform similarly to standard BN.
BN Method $\epsilon=0.03$ $\epsilon=0.1$ $\epsilon=0.15$ $\epsilon=0.2$ $\epsilon=0.3$
Standard (32) 0.26 $\pm$ 0.04 0.07 $\pm$ 0.05 0.18 $\pm$ 0.18 0.45 $\pm$ 0.53 1.0 $\pm$ 0.96
[…]
Standard (128) 0.05 $\pm$ 0.02 0.06 $\pm$ 0.05 0.05 $\pm$ 0.02 0.11 $\pm$ 0.04 0.39 $\pm$ 0.13
Ours (32) 42.72 $\pm$ 1.16 26.06 $\pm$ 1.5 21.66 $\pm$ 1.31 18.74 $\pm$ 0.99 14.868 $\pm$ 0.94
Ours (64) 43.29 $\pm$ 0.56 26.17 $\pm$ 0.41 21.66 $\pm$ 0.34 18.69 $\pm$ 0.39 14.8 $\pm$ 0.35
Ours (128) 36.36 $\pm$ 1.38 20.27 $\pm$ 1.35 16.82 $\pm$ 1.27 14.41 $\pm$ 1.16 11.45 $\pm$ 0.97
Mean corrected (32) 34.39 $\pm$ 0.84 20.07 $\pm$ 0.72 16.53 $\pm$ 0.76 14.28 $\pm$ 0.6 11.42 $\pm$ 0.49
Mean corrected (64) 26.82 $\pm$ 1.08 14.4 $\pm$ 1.16 11.99 $\pm$ 0.99 10.29 $\pm$ 0.84 7.98 $\pm$ 0.59
Mean corrected (128) 16.57 $\pm$ 0.96 8.43 $\pm$ 1.19 7.03 $\pm$ 1.17 6.14 $\pm$ 1.0 4.92 $\pm$ 0.83
Khoshsirat et al. (32) 37.04 $\pm$ 0.71 22.78 $\pm$ 0.92 19.06 $\pm$ 0.86 16.35 $\pm$ 0.66 12.75 $\pm$ 0.49
Khoshsirat et al. (64) 38.27 $\pm$ 1.01 22.97 $\pm$ 0.49 19.07 $\pm$ 0.62 16.24 $\pm$ 0.73 12.3 $\pm$ 0.64
Khoshsirat et al. (128) 33.45 $\pm$ 0.79 18.49 $\pm$ 0.82 15.22 $\pm$ 0.71 12.94 $\pm$ 0.54 9.9 $\pm$ 0.85
Table 4: Performance of ResNet-9 with different BN methods under varying FGSM attack and batch sizes respectively. Values depict mean accuracy $\pm$ standard deviation on test CIFAR-10 data. Numbers inside the parentheses in the far left column denote batch size.
BN Method $\epsilon=0.03$ $\epsilon=0.1$ $\epsilon=0.15$ $\epsilon=0.2$ $\epsilon=0.3$
Standard (8) 9.4 $\pm$ 1.5 7.6 $\pm$ 0.9 6.2 $\pm$ 1.7 4.9 $\pm$ 2.3 3.8 $\pm$ 2.6
Standard (16) 10.2 $\pm$ 1.0 5.5 $\pm$ 1.5 4.7 $\pm$ 1.7 4.3 $\pm$ 1.9 3.9 $\pm$ 2.0
Ours (8) 38.1 $\pm$ 1.4 28.8 $\pm$ 1.8 25.9 $\pm$ 2.4 23.6 $\pm$ 3.2 20.9 $\pm$ 4.5
Ours (16) 35.4 $\pm$ 0.6 24.9 $\pm$ 1.0 21.8 $\pm$ 1.3 19.8 $\pm$ 1.6 17.3 $\pm$ 1.8
Mean corrected (8) 33.7 $\pm$ 1.9 24.4 $\pm$ 1.4 22.3 $\pm$ 1.2 20.8 $\pm$ 1.3 18.9 $\pm$ 1.8
Mean corrected (16) 33.5 $\pm$ 2.7 24.6 $\pm$ 3.3 21 $\pm$ 3.1 19.4 $\pm$ 2.9 17 $\pm$ 3.3
Khoshsirat et al. (8) 35.9 $\pm$ 2.2 26.2 $\pm$ 2.6 23.4 $\pm$ 2.6 21.7 $\pm$ 2.6 19.7 $\pm$ 2.4
Khoshsirat et al. (16) 33.4 $\pm$ 1.8 22.9 $\pm$ 2.3 20.4 $\pm$ 2.4 18.87 $\pm$ 2.3 16.9 $\pm$ 2.1
Lasso (8) 10.5 $\pm$ 5.7 5.7 $\pm$ 3.4 5.0 $\pm$ 2.9 4.8 $\pm$ 2.7 4.8 $\pm$ 2.8
Lasso (16) 6.8 $\pm$ 2.9 3.6 $\pm$ 1.8 3.3 $\pm$ 1.7 3.5 $\pm$ 1.7 4.2 $\pm$ 2.3
Ridge (8) 9.7 $\pm$ 1.8 6.2 $\pm$ 0.8 4.8 $\pm$ 0.8 4.3 $\pm$ 0.7 4.1 $\pm$ 0.8
Ridge (16) 12.9 $\pm$ 6.3 7.0 $\pm$ 1.0 5.9 $\pm$ 0.9 5.5 $\pm$ 0.9 4.9 $\pm$ 1.0
Table 5: Performance of HRNetV2 with different BN methods under varying FGSM attack and batch sizes respectively. Values depict mean of mIoU scores $\pm$ standard deviation on part of validation Cityscapes data using as a test data. Numbers inside the parentheses in the far left column denote batch size.
In addition, we evaluated robustness under the PGD attack. The experimental configuration mirrors that of the FGSM case; The corresponding results are summarized in Tables 6, 7.
BN Method $\epsilon=0.03$ $\epsilon=0.1$ $\epsilon=0.15$ $\epsilon=0.2$ $\epsilon=0.3$
Standard (32) 0.08 $\pm$ 0.02 0.08 $\pm$ 0.01 0.0 0.0 0.0
Standard (64) 0.19 $\pm$ 0.03 0.0 0.0 0.0 0.0
Standard (128) 0.0 0.0 0.0 0.0 0.0
Ours (32) 27.46 $\pm$ 0.74 17.27 $\pm$ 0.48 13.5 $\pm$ 0.81 10.74 $\pm$ 0.64 6.8 $\pm$ 0.47
Ours (64) 32.05 $\pm$ 1.86 21.7 $\pm$ 1.5 17.01 $\pm$ 1.4 13.31 $\pm$ 1.18 8.44 $\pm$ 0.69
Ours (128) 24.54 $\pm$ 1.7 14.86 $\pm$ 1.54 11.29 $\pm$ 1.39 8.88 $\pm$ 1.08 5.6 $\pm$ 0.5
Mean corrected (32) 19.57 $\pm$ 1.61 10.84 $\pm$ 1.73 7.83 $\pm$ 1.67 5.98 $\pm$ 1.38 3.52 $\pm$ 0.94
Mean corrected (64) 13.65 $\pm$ 1.93 7.03 $\pm$ 1.93 4.96 $\pm$ 1.46 3.7 $\pm$ 1.24 2.01 $\pm$ 0.81
Mean corrected (128) 5.3 $\pm$ 0.71 2.05 $\pm$ 0.38 1.45 $\pm$ 0.27 0.94 $\pm$ 0.23 0.51 $\pm$ 0.11
Khoshsirat et al. (32) 28.36 $\pm$ 1.44 20.302 $\pm$ 1.6 14.61 $\pm$ 1.27 10.37 $\pm$ 1.15 5.3 $\pm$ 0.6
Khoshsirat et al. (64) 28.62 $\pm$ 1.08 20.45 $\pm$ 1.02 14.76 $\pm$ 1.15 10.37 $\pm$ 0.64 5.25 $\pm$ 0.42
Khoshsirat et al. (128) 25.5 $\pm$ 0.47 17.74 $\pm$ 0.54 12.53 $\pm$ 0.04 7.83 $\pm$ 0.14 3.95 $\pm$ 0.2
Table 6: Performance of ResNet-9 with different BN methods under varying PGD attack and batch sizes respectively. Values depict mean accuracy $\pm$ standard deviation on test CIFAR-10 data. Numbers inside the parentheses in the far left column denote batch size.
BN Method $\epsilon=0.03$ $\epsilon=0.1$ $\epsilon=0.15$ $\epsilon=0.2$ $\epsilon=0.3$
Standard (8) 1.8 $\pm$ 1.4 0.8 $\pm$ 0.7 0.6 $\pm$ 0.5 0.5 $\pm$ 0.3 0.6 $\pm$ 0.4
Standard (16) 2.0 $\pm$ 1.0 0.5 $\pm$ 0.2 0.3 $\pm$ 0.1 0.2 $\pm$ 0.1 0.2 $\pm$ 0.0
Ours (8) 10.4 $\pm$ 1.1 4.7 $\pm$ 1.1 3.4 $\pm$ 1.1 2.7 $\pm$ 1.1 2.2 $\pm$ 1.1
Ours (16) 10.6 $\pm$ 0.9 5.0 $\pm$ 0.8 3.7 $\pm$ 0.8 2.9 $\pm$ 0.8 2.2 $\pm$ 0.6
Mean corrected (8) 10.1 $\pm$ 0.9 3.7 $\pm$ 0.6 2.4 $\pm$ 0.5 1.8 $\pm$ 0.4 1.3 $\pm$ 0.4
Mean corrected (16) 10.5 $\pm$ 1.9 4.8 $\pm$ 1.3 3.4 $\pm$ 1.1 2.7 $\pm$ 0.9 2.0 $\pm$ 0.7
Khoshsirat et al. (8) 8.5 $\pm$ 1.0 2.6 $\pm$ 0.2 1.7 $\pm$ 0.2 1.4 $\pm$ 0.2 1.1 $\pm$ 0.2
Khoshsirat et al. (16) 9.1 $\pm$ 1.6 3.3 $\pm$ 1.0 2.3 $\pm$ 0.6 1.8 $\pm$ 0.4 1.4 $\pm$ 0.2
Lasso (8) 2.8 $\pm$ 1.4 2.0 $\pm$ 1.8 1.9 $\pm$ 2.0 1.8 $\pm$ 2.2 1.6 $\pm$ 2.3
Lasso (16) 2.6 $\pm$ 1.5 1.6 $\pm$ 1.0 1.4 $\pm$ 0.9 1.3 $\pm$ 1.0 1.1 $\pm$ 1.0
Ridge (8) 2.6 $\pm$ 1.8 1.4 $\pm$ 1.8 1.1 $\pm$ 1.6 0.9 $\pm$ 1.2 0.5 $\pm$ 0.5
Ridge (16) 2.8 $\pm$ 0.6 1.0 $\pm$ 0.4 0.9 $\pm$ 0.6 0.8 $\pm$ 0.7 0.7 $\pm$ 0.8
Table 7: Performance of HRNetV2 with different BN methods under varying PGD attack and batch sizes respectively. Values depict mean of mIoU scores $\pm$ standard deviation on part of validation Cityscapes data using as a test data. Numbers inside the parentheses in the far left column denote batch size.
Overall, the findings support our theoretical results: Applying JS-based shrinkage to both the mean and variance yields significantly improved robustness under adversarial settings compared to other BN methods.
5 Discussion and Conclusion
One possible explanation for the improved performance of JS-BN under adversarial perturbations lies in its effect on the Lipschitz properties of the network. It is well known that smaller local Lipschitz constants are associated with more regular functions, which are inherently more stable to perturbations (Evans and Gariepy, 2015). From the empirical side, recent work has shown that local Lipschitzness enforces both robustness and accuracy (Yang et al., 2020). Our theoretical results show that JS-BN reduces the local Lipschitz constant relative to standard BN. Intuitively, this shrinkage effect smooths the geometry of the loss landscape.
For BN, we have the following.
$f(x)=\gamma\frac{x-\mu}{\sqrt{\sigma^{2}+\epsilon}}+\beta,$
which yields
$|f(x)-f(y)|=\left|\frac{\gamma}{\sqrt{\sigma^{2}+\epsilon}}(x-y)\right|.$
Thus, the corresponding Lipschitz constant is
$L=\frac{|\gamma|}{\sqrt{\sigma^{2}+\epsilon}}.$
Since, by construction, the JS-shrinkage produces a larger variance estimate, $\sigma^{2,\,\mathrm{JS}}>\sigma^{2}$, the denominator increases, resulting in a smaller local Lipschitz constant. This naturally leads to an improved smoothness and robustness of the model.
In summary, in this paper, we present a novel theory that demonstrates the dominance of Stein’s shrinkage estimators for mean and variance over SM and SV under adversarial attacks. We applied JS-BN in a variety of DNNs and demonstrated superior performance compared to other BN methods on three benchmark datasets. By construction, Stein’s shrinkage improves Lipschitzness and therefore robustness of the model, while maintaining high accuracy.
Acknowledgments and Disclosure of Funding